# Engagement Maximization

### v4

@@ 2.3 Discussion
The generalized entropy functions $G$ and $H$ play conceptually distinct roles in our model. The function $H$ governs the agent’s information processing constraint. It determines which kinds of information are relatively easy or difficult for the agent to process. The function $G$, on the other hand, measures the kind of engagement that generates profits for the principal. One leading case of interest is when $H=G$. We interpret this case as one in which the principal displays advertisements to the user interspersed with content, and the principal is paid “per impression” by advertisers without any direct measurement of whether the user paid attention to the advertisement. The principal’s profits are maximized by maximizing the time the agent spends attending to the principal’s content. Arrangements similar to this assumption are observed in a variety of media markets (see, e.g., Gentzkow et al. (2024)). [^10] The authors would like to thank Emir Kamenica for suggesting this interpretation. Note that when $H=G$, the principal’s objective is not to maximize the stopping time $\tau$ directly; however, in equilibrium the constraint (2) will bind, and the principal will maximize time spent spend subject to (2) holding with equality.
A second leading case is one in which $G$ measures advertisement-related information acquisition. This information can be partially or completely irrelevant to the agent’s decision problem, which is to say that there are beliefs $q,q^{\prime}\in\mathcal{P}(X)$ that are identical from the agent’s perspective ($\hat{u}(q)=\hat{u}(q^{\prime})$) but distinct from the principal’s perspective ($G(q)\neq G(q^{\prime})$). We interpret this case as one which the principal cares about whether or not the agent pays attention to the advertisements that are embedded in the content the principal provides, perhaps because part of the principal’s profits (now or in the future) are based on the effectiveness of the advertisements. Section 5.6 offers examples of these two cases, in a different interpretative context (engaging test-motivated students). Note also that these two cases are not mutually exclusive; a principal could be paid per impression at a rate that is linked to the historical ad effectiveness, which could be modeled in our framework as setting $G$ equal to the sum of $H$ and another entropy function measuring the attention paid to advertisements.
@@
Our formulation of the principal’s problem (Definition 1) embeds several assumptions. First, we assume that the principal acts first and can commit to law $P$ before the agent chooses her stopping strategy $\tau$. Given that the agent faces a pre-determined $P$, her problem is dynamically consistent and the agent’s commitment or lack thereof is irrelevant. In the context of an internet platform and a user, we interpret this assumption as implying that the internet platform designs its recommendation algorithms before the user can commit to a usage pattern. We discuss the role of commitment for the principal in Section 4.
@@ 3.1 A Relaxed Problem
Lemma 1 presents a necessary condition for any admissible policy for the principal in the optimization problem (Definition 1). The first condition states that the agent’s optimal stopping utility is weakly greater than the utility from stopping immediately. The second condition states that the cumulative information acquired by the agent is weakly less than $\chi\mathbb{E}[\tau|\mathcal{F}_{0}]$—the maximal attention permitted by the information constraint (2).
@@
$\bar{J}(q_{0})=\sup_{\begin{subarray}{c}\pi\in\mathcal{P}(\mathcal{P}(X)):\mathbb{E}^{\pi}[q]=q_{0}\end{subarray}}$ $\mathbb{E}^{\pi}[G(q)-G(\bar{q}_{0})]$ (3)
@@
Equation 3 is a convex optimization problem that satisfies the conditions in Theorem 4 of Zhong (2018), which immediately implies:
@@
Equation 3 has a solution $\pi^{*}$ with finite support. Either $\bar{J}(q_{0})=0$ or $\bar{J}(q_{0})>0$ and there exists $\lambda>0$ s.t. all $\pi^{*}$ satisfy
@@
In both cases, $\frac{\bar{\kappa}}{\chi}\mathbb{E}^{\pi^{*}}[H(q)-H(\bar{q}_{0})]=\mathbb{E}^{\pi^{*}}[\widehat{\mu}(q)-\widehat{\mu}(\bar{q}_{0})]$.
@@ 3.2 Implementation
For all non-degenerate $\pi\in\mathcal{P}(\mathcal{P}(X))$ that satisfy the constraint in (3), let $\alpha=\frac{\chi}{\mathbb{E}^{\pi}[H(q)-H(\bar{q}_{0})]}$ and let $q_{t}$ be the $\alpha$-dilution of $\pi$. Then $q_{t}$ is feasible in the problem of Definition 1 and implements utility level $\mathbb{E}^{\pi}[G(q)-G(\bar{q}_{0})]$.
@@
Take any admissible stopping time $\tau^{\prime}$,
$=$ $\mathbb{E}^{P}[\widehat{\mu}(q)-\bar{\kappa}\tau|\mathcal{F}_{0}].$
The first equality is from the definition of conditional expectations and the process defining $q_{t}$, (5). The first inequality is from the constraint $\mathbb{E}^{\pi}[\widehat{\mu}(q)-\widehat{\mu}(\bar{q}_{0})]\geq\bar{\kappa}\mathbb{E}^{P}[\tau|\mathcal{F}_{0}]$ in the relaxed problem (3) and $E[\tau|\tau^{\prime}\geq\tau]\leq E[\tau^{\prime}|\tau^{\prime}\geq\tau]$. The second equality is from the memorylessness property of $\tau$, [^14] Specifically, $\mathbb{E}^{P}\left[\tau|\tau^{\prime}<\tau\right]=\mathbb{E}^{P}\left[\tau-\tau^{\prime}|\tau^{\prime}<\tau\right]+\mathbb{E}^{P}\left[\tau^{\prime}|\tau^{\prime}<\tau\right]=\mathbb{E}^{P}\left[\tau|\mathcal{F}_{0}\right]+\mathbb{E}^{P}\left[\tau^{\prime}|\tau^{\prime}<\tau\right]$. and the last from the definition of conditional expectations. Note that if the constraint in (3) is slack, the first inequality is strict if $\tau^{\prime}$ is not equal to $\tau$ $P$-a.e. In this case, the $\alpha$-dilution of $\pi$ strongly implements principal’s utility level $\rho\mathbb{E}^{\pi}[G(q_{t})-G(\bar{q}_{0})]$. ∎
@@
$\forall\bar{q}_{0}\in\mathcal{P}(X)$, there exists $\pi^{*}\in\mathcal{P}(\mathcal{P}(X))$ with finite support solving Equation 3. If Supp$(\pi^{*})=\{\bar{q}_{0}\}$, the agent will immediately stop and any feasible policy is optimal. Otherwise, let $\alpha^{*}=\frac{\chi}{\mathbb{E}^{\pi^{*}}[H(q)-H(\bar{q}_{0})]}$, and let $(P^{*},\tau^{*})$ be the law and jumping time of the $\alpha^{*}$-dilution of $\pi^{*}$. Then, $(P^{*},I_{t}^{H*}=\chi t,\tau^{*})$ is an optimal solution to the principal’s problem (Definition 1).
@@
There are generally many optimal policies in the principal’s problem. First, there may be multiple $\pi^{*}$ that solve the relaxed principal’s problem (although uniqueness in static rational inattention problems is guaranteed under certain additional assumptions). Second, there are many stochastic processes $q_{t}$ (equivalently, laws $P$) and stopping times $\tau$ that induce the same law for $q_{\tau}$; provided that this law is equal to $\pi^{*}$ and that incentive compatibility and the bounds on information acquisition are satisfied, all such policies are optimal. However, it is not the case that anything goes, as we show next.
3.3 Characterization of optimal policies
For the purpose of discussing optimal polices, we assume for this subsection that the optimal policy is to acquire some information ($\bar{J}(\bar{q}_{0})>0$ in the context of Proposition 1).
To begin, we define, given a measure $\pi\in\mathcal{P}(\mathcal{P}(X))$, the value of information acquisition in a hypothetical restricted static rational inattention problem:
$V_{R}(q,\pi)=\max_{\pi^{\prime}\in\mathcal{P}(\text{Supp}(\pi)):E^{\pi^{\prime}}[q^{\prime}]=q}\mathbb{E}^{\pi^{\prime}}\left[\widehat{\mu}(q^{\prime})-\hat{u}(q)-\frac{\bar{\kappa}}{\chi}(H(q^{\prime})-H(q))\right].$
This problem maximizes the agent’s expected utility subject to the usual constraint of Bayes consistency and the additional constraint that the posteriors lie in the support of $\pi$. We use the notation $V_{R}(q,\pi)$ to emphasize that the problem is restricted relative to the usual rational inattention problem. Note that the problem is feasible only for $q$ that lie in in the convex hull of the support of $\pi$, which we denote by $\text{Conv}(\text{Supp}(\pi))$.
We interpret $V_{R}(q,\pi)$ as describing the value of information acquisition, in the sense that $V_{R}(q,\pi)$ is the difference between the utility the agent achieves with and without information acquisition. [^15] The value of information acquisition is related to, but not the same as, the notion of uncertainty defined in Frankel and Kamenica (2019). We define the set of “suspensive” beliefs given $\pi$ as the set for which the value of information acquisition is weakly positive.
Definition 2.
Given $\pi\in\mathcal{P}(\mathcal{P}(X))$, the belief $q\in\mathcal{P}(X)$ is decisive if $q\in\text{Supp}(\pi)$; the belief $q\in\mathcal{P}(X)$ is suspensive if $q\text{$\in$Conv}(\text{Supp}(\pi))\setminus\mathrm{Supp(\pi)}$ and satisfies $V_{R}(q,\pi)\geq 0$. Let $\Delta(\pi)\subseteq P(X)$ denote the set of all suspensive beliefs given $\pi$.
Given a measure $\pi$ that describes the stopping beliefs, it is intuitive that beliefs in $\mathrm{Supp(\pi)}$ are “decisive” because they directly lead to stopping. At the suspensive beliefs in $\Delta(\pi)$, the agent benefits, relative to the prior consistent with $\pi$, from the signals that the principal is offering (which always result in some posterior in the support of $\pi$). Therefore, $\Delta(\pi)$ characterizes the beliefs at which the principal can possibly induce “suspense” before eventually leading the agent to stopping beliefs in $\mathrm{Supp(\pi)}$.
Suppose the optimal policy in (3), $\pi^{*}$, is unique. In this case, the stochastic process for beliefs $q_{t}$ cannot leave the set of suspensive beliefs $\Delta(\pi^{*})$ before $t=\tau$. If $q_{t}$ left the convex hull of the support of $\pi^{*}$, then $\pi^{*}$ would not describe the law of $q_{\tau}$. If $q_{t}$ was decisive or if $V_{R}(q_{t},\pi^{*})<0$, the agent would choose to stop. In the case of multiple optimal policies, this argument applies for some optimal $\pi^{*}$, which leads to the following result.
Proposition 3.
Given $\bar{q}_{0}\in\mathcal{P}(X)$, let $\Pi^{*}$ be the set of solutions to (3). Then, in all solutions to the principal’s problem, for all $t\in[0,\tau^{*})$, beliefs are suspensive given some $\pi^{*}\in\Pi^{*}$,
$\text{Supp}(q_{t})\subseteq\bigcup_{\pi^{*}\in\Pi^{*}}\Delta(\pi^{*}),$
and at $t=\tau^{*}$, beliefs are decisive given some $\pi^{*}\in\Pi^{*}$,
$\text{Supp}(q_{\tau})\subseteq\bigcup_{\pi^{*}\in\Pi^{*}}\text{Supp}(\pi^{*}).$
Proof.
Let $\pi^{*}$ be a solution to the problem defined in (3). We first show that for any $q^{\prime}\in\text{Supp}(\pi^{*})$, $\bar{J}(q^{\prime})=0.$ Let $\pi^{**}$ be the optimal policy associated with $q^{\prime}$. The policy $\pi^{***}(q)=\pi^{*}(q)\mathbf{1}\{q\neq q^{\prime}\}+\pi^{*}(q^{\prime})\pi^{**}(q)$ is feasible in (3) given the prior $\bar{q}_{0}$ and would achieve strictly higher utility than $\pi^{*}$ if $\bar{J}(q^{\prime})>0$. It follows in the principal-agent problem that if $q_{t}\in\text{Supp}(\pi^{*})$, then $t=\tau^{*}$, because there is no way to induce the agent to continue.
By Theorem 1, there is some $\pi^{*}\in\Pi^{*}$ that is the law of $q_{\tau^{*}}$ under the optimal policies, and the second claim follows. By the martingale property of the belief process, $q_{t}=E^{P^{*}}[q_{\tau}|\mathcal{F}_{t},t<\tau^{*}]$ for all $t\in[0,\tau^{*})$, and thus $q_{t}$ must lie in the convex hull of the support of $\pi^{*}$. The belief $q_{t}$ cannot be decisive if $t<\tau^{*}$, by the argument above. By the definition of (3), $\frac{\bar{\kappa}}{\chi}\mathbb{E}^{\pi^{*}}[H(q)-H(\bar{q}_{0})]\leq\mathbb{E}^{\pi^{*}}[\widehat{\mu}(q)-\widehat{\mu}(\bar{q}_{0})]$, from which it follows immediately that $V_{R}(q_{t},\pi^{*})\geq 0$. ∎
Solving (3) generically leads to the IC constraint binding, which suggests that the value of acquiring $\pi^{*}$ is exactly zero at the prior for the agent. The set of suspensive beliefs $\Delta(\pi^{*})$ thus defines beliefs that are “more uncertain” than the prior. Therefore, the interpretation of Proposition 3 is that two types of signals appear in any optimal policy almost surely. The first type are decisive signals that are so informative that the agent makes a decision immediately after receiving the signal. The second type are suspensive signals that causes the agent’s beliefs to move to a posterior that is “more uncertain” than the prior.
3.4 The agent-optimal benchmark
@@
This solution can be implemented, as above, by a compound Poisson process, but also by a diffusion (and many other processes).
@@
This static rational inattention problem is essentially identical to the one that characterizes optimal policy in our principal-agent framework, except that the UPS information cost is $\frac{\bar{\kappa}}{\chi}H(q)$ in the benchmark problem and $\frac{\bar{\kappa}}{\chi}H(q)-\lambda G(q)$ in the principal agent case. The similarity in structure between these two problems will allow us to highlight the implications of engagement maximization.
Before proceeding, we should note that the principal is unable to induce any information acquisition whenever the only feasible $\pi$ in the relaxed problem places full support on $\bar{q}_{0}$; this corresponds to the case in the benchmark model in which the agent strictly prefers to not gather information. In all other cases, the principal can induce information acquisition in the principal-agent model.
3.5 Example (continued)
We illustrate Theorem 1 by continuing the numerical example in Section 1.1. Note in this example that $H=G$ (both are the negative of Shannon’s entropy). We maintain the same model primitives, with the exceptions that we now vary the agent’s prior belief and focus our analysis on the agent’s posteriors when stopping ($q_{\tau}$) as opposed to the compound Poisson signals that generate those posteriors. Figure 1 illustrates the optimal $q_{\tau}$ for every prior belief: When $\bar{q}_{0}\leq\underline{q}$ or $\bar{q}_{0}\geq\bar{q}$, the agent stops immediately no matter what information she will receive. When $\bar{q}_{0}\in(\underline{q},\bar{q})$, the dashed lines illustrate the support of the agent-optimal policy. The lines are flat, meaning that the agent’s stopping beliefs do not change with the prior belief, consistent with the prediction of the standard RI models. [^16] This property is known as the “locally invariant posteriors” (Caplin et al. (2022)). The solid curves illustrate the support of the unique optimal $\pi^{*}$. They are closer to the boundaries zero and one, indicating that the posterior beliefs become more polarized under the principal’s optimal policy.
Figure 1: $\text{Supp}(q_{\tau})$ as a correspondence of $q_{0}$
Figure 2: Engagement as a function of $q_{0}$
Now, focus on the $\bar{q}_{0}=\frac{1}{2}$ case. As is illustrated in Figure 1, the unique optimal $\pi^{*}$ involves two posterior beliefs $\left\{q^{1},q^{2}\right\}$, where $q^{1}<\underline{q}$ and $q^{2}>\bar{q}$. Per Proposition 2, the dilution of $\pi^{*}$ implements an optimal policy. In this case, the dilution of $\pi^{*}$ is the unique optimal policy. It is apparent from the symmetry of the problem that the value of information acquisition, $V_{R}(q,\pi^{*})$, is maximized at $q=\frac{1}{2}$. As a result, the set of suspensive beliefs, $\Delta(\pi^{*})$, is a singleton, $\{\bar{q}_{0}\}$. By Proposition 3, beliefs will remain at $\bar{q}_{0}$ until they jump to either $q^{1}$ or $q^{2}$.
When $\bar{q}_{0}<\frac{1}{2}$, the optimal policy is not unique. In this case, again by the symmetry of the problem, it is apparent that the set of suspensive beliefs is $\Delta(\pi^{*})=[q_{0},1-q_{0}]$. Beliefs will remain in this interval until eventually jumping to either $q^{1}$ or $q^{2}$. A symmetric argument applies when $\bar{q}_{0}>\frac{1}{2}$. Figure 3 illustrates the sample paths of one optimal policy, in which beliefs jump between $\bar{q}_{0}$ and $1-\bar{q}_{0}$ before eventually jumping to $q^{1}$ or $q^{2}$.
Figure 3: Sample paths of an optimal policy
4 Implications
In this section, we consider several implications of our main result.

### v5

@@ 2.3 Discussion
The generalized entropy functions $G$ and $H$ play conceptually distinct roles in our model. The function $H$ governs the agent’s information processing constraint. It determines which kinds of information are relatively easy or difficult for the agent to process. The function $G$, on the other hand, measures the kind of engagement that generates profits for the principal. One leading case of interest is when $H=G$.
@@
In the context of an internet platform, we interpret this case as one in which the principal displays advertisements to the user interspersed with content, and the principal is paid “per impression” by advertisers without any direct measurement of whether the user paid attention to the advertisement. The principal’s profits are maximized by maximizing the time the agent spends attending to the principal’s content. Arrangements similar to this assumption are observed in a variety of media markets (see, e.g., Gentzkow et al. (2024)). [^10] The authors would like to thank Emir Kamenica for suggesting this interpretation. Note that when $H=G$, the principal’s objective is not to maximize the stopping time $\tau$ directly; however, in equilibrium the constraint (2) will bind, and the principal will maximize time spent spend subject to (2) holding with equality. In the context of a teacher attempting to engage a test-motivated student, we interpret the $H=G$ case as one in which the teacher’s goal is for the student to learn as much as possible, without any particular preference for whether the student learns test-relevant or test-irrelevant information.
A second leading case is one in which $G$ measures only information acquisition that is irrelevant to the agent’s decision problem. In the context of an internet platform, we interpret this case as one in which the platform profits from the users’ attention to advertisements that are unrelated to the content the user is interested in. This case differs from the $H=G$ case in that the internet platform profits only if the user actually pays attention to the advertisement. This would be the case if there is some measurement of the effectiveness of the ad (using clicks, cookies, or other trackers), and the platform earns revenue on the basis of this measurement. In the context of engaging test-motivated students, we interpret this case as one in which the teacher cares only about whether the student learns the course material and not about whether the student does well on the test, which is the exact opposite of what the student cares about.
In both the internet platform and teacher-student context, it is natural to assume that elements of both of these two leading cases are present. An internet platform will profit on both a per-impression basis and on the basis of whether ads on its platform are effective, which can be modeled in our framework as setting $G$ equal to the sum of $H$ and another entropy function measuring the attention paid to advertisements. This second component of $G$ could be irrelevant to the agent’s decision problem, if the ads are in fact unrelated, or not, if the ads are related to information the user is attempting to acquire. Likewise, in the teacher-student context, a teacher might care both about whether the student passes the test and whether the student learns the course material (including content that can and cannot be tested in an exam). This can be modeled in our framework as setting $G$ equal to the sum of $\hat{u}$ (representing the student’s expected test performance) and a function measuring the student’s understanding of the course content. [^11] Our preferred interpretation of the model is that the principal cares only about the information the agent acquires and not the action the agent selects. However, this example shows that our model also covers the case in which the principal and agent agree on the optimal action conditional on the agent’s beliefs (although our extension to the case of an informed principal in Section 5.2 would not hold in this case). Our assumption that $A$ is finite implies that $\hat{u}$ is not continuously differentiable, and hence this example would require minor modifications of some of our technical assumptions.
@@
Apart from the distinction between $H$ and $G$, our formulation of the principal’s problem (Definition 1) embeds several assumptions. First, we assume that the principal acts first and can commit to law $P$ before the agent chooses her stopping strategy $\tau$. Given that the agent faces a pre-determined $P$, her problem is dynamically consistent and the agent’s commitment or lack thereof is irrelevant. In the context of an internet platform and a user, this assumption can be motivated by assuming that the internet platform designs its recommendation algorithms before the user can commit to a usage pattern. However, this justification is inconsistent with the platform’s ability to change its algorithms at any time without informing its users, and is not particularly appealing in the teacher-student context either. For this reason, we will devote a significant part of our analysis to the case without commitment in Section 4.2, and demonstrate that the solutions we construct in the commitment case can also be used to characterize equilibria in the case without commitment.
@@ 3.1 A Relaxed Problem
Lemma 1 presents a necessary condition for any admissible policy for the principal in the optimization problem (Definition 1). The first condition states that the agent’s optimal stopping utility is weakly greater than the utility from stopping immediately. The second condition states that the expected cumulative information acquired by the agent is weakly less than $\chi\mathbb{E}[\tau|\mathcal{F}_{0}]$—the maximal attention permitted by the information constraint (2).
@@
$\bar{J}(q_{0})=\sup_{\begin{subarray}{c}\pi\in\mathcal{P}(\mathcal{P}(X)):\mathbb{E}^{\pi}[q]=q_{0}\end{subarray}}$ $\mathbb{E}^{\pi}[G(q)-G(\bar{q}_{0})]$ (R)
@@
Equation R has a solution $\pi^{*}$ with finite support. Either $\bar{J}(q_{0})=0$ or $\bar{J}(q_{0})>0$ and there exists $\lambda>0$ s.t. all $\pi^{*}$ satisfy
@@
In both cases, $\frac{\bar{\kappa}}{\chi}\mathbb{E}^{\pi^{*}}[H(q)-H(\bar{q}_{0})]=\mathbb{E}^{\pi^{*}}[\widehat{\mu}(q)-\widehat{\mu}(\bar{q}_{0})]$ and $\mathbb{E}^{\pi^{*}}[\bar{J}(q)]=0$.
Proof.
See Appendix A.1. ∎
@@ 3.2 Implementation
For all non-degenerate $\pi\in\mathcal{P}(\mathcal{P}(X))$ that satisfy the constraint in (R), let $\alpha=\frac{\chi}{\mathbb{E}^{\pi}[H(q)-H(\bar{q}_{0})]}$ and let $q_{t}$ be the $\alpha$-dilution of $\pi$. Then $q_{t}$ is feasible in the problem of Definition 1 and implements utility level $\mathbb{E}^{\pi}[G(q)-G(\bar{q}_{0})]$.
@@
See Appendix A.2. ∎
@@
$\forall\bar{q}_{0}\in\mathcal{P}(X)$, there exists $\pi^{*}\in\mathcal{P}(\mathcal{P}(X))$ with finite support solving Equation R. If Supp$(\pi^{*})=\{\bar{q}_{0}\}$, the agent will immediately stop and any feasible policy is optimal. Otherwise, let $\alpha^{*}=\frac{\chi}{\mathbb{E}^{\pi^{*}}[H(q)-H(\bar{q}_{0})]}$, and let $(P^{*},\tau^{*})$ be the law and jumping time of the $\alpha^{*}$-dilution of $\pi^{*}$. Then, $(P^{*},I_{t}^{H*}=\chi t,\tau^{*})$ is an optimal solution to the principal’s problem (Definition 1).
@@
There are generally many optimal policies in the principal’s problem. First, there may be multiple $\pi^{*}$ that solve the relaxed principal’s problem (although uniqueness in static rational inattention problems is guaranteed under certain additional assumptions). Second, there are many stochastic processes $q_{t}$ (equivalently, laws $P$) and stopping times $\tau$ that induce the same law for $q_{\tau}$; provided that this law is equal to $\pi^{*}$ and that incentive compatibility and the bounds on information acquisition are satisfied, all such policies are optimal. However, it is not the case that anything goes, as we show in section 4.
3.3 The agent-optimal benchmark and belief polarization
@@
This solution can be implemented, as above, by a compound Poisson process, but also by a diffusion (and many other processes). The following characterization is straightforward.
Lemma 2.
$V^{B}(q)>\hat{u}(q)$implies $\bar{J}(q)>0$. Moreover, $\forall\pi$ that implements $V^{B}(\bar{q}_{0})$, $\mathbb{E}^{\pi}[V^{B}(q)-\hat{\mu}(q)]=0$.
Proof.
See Appendix A.3. ∎
This lemma highlights two points. First, if at some $q$ the agent would choose to acquire information in the agent-optimal benchmark, the principal can profit by providing the agent with information (although that information will not in general be the information the agent would have chosen to acquire). Second, at any belief the agent might hold when stopping in the agent-optimal benchmark, the agent must have no benefit from continuing to acquire information.
@@
The static rational inattention problem that characterizes stopping beliefs in the agent-optimal benchmark is essentially identical to the one that characterizes optimal policy in our principal-agent framework, except that the UPS information cost is $\frac{\bar{\kappa}}{\chi}H(q)$ in the benchmark problem and $\frac{\bar{\kappa}}{\chi}H(q)-\lambda G(q)$ in the principal agent case. The similarity in structure between these two problems will allow us to highlight the implications of engagement maximization.
Proposition 3.
Proof.
See the appendix, section A.4. ∎
That is, the stopping beliefs in the principal-agent problem will lie outside the convex hull of the stopping beliefs of the agent-optimal benchmark. When $G$ is strictly convex, it is also possible to show that the stopping beliefs in the agent-optimal benchmark are in the relative interior of the convex hull of the stopping beliefs in the principal-agent problem. However, when the principal profits only by providing the agent with irrelevant information, $G$ is not strictly convex, and the agent-optimal benchmark stopping beliefs will not necessarily lie in the relative interior. Figure 4 below provides an example of this possibility.
Leading Case: $H$=$G$.
When the principal benefits from all forms of engagement by the agent (our interpretation of the $H=G$ case), it is never worthwhile for the principal to provide the agent with decision-irrelevant information. Instead, the principal induces the agent to acquire more decision-relevant information than the agent would choose to acquire on her own.
To illustrate this, we suppose that the state space has a product structure, $X=X_{1}\times X_{2}$, and that the agent’s decision problem depends only on $x_{1}\in X_{1}$. We further suppose that $x_{1}$ and $x_{2}$ are independent under the prior, so that an agent who receives signals that depend only on $x_{1}$ will not update about $x_{2}$.
The following result shows that this is what happens in the principal-agent problem– the principal will send the agent information about $x_{1}$ only. This occurs even though the principal would benefit from the agent learning about $x_{2}$ as well. Intuitively, signals about $x_{1}$ and $x_{2}$ that benefit the principal equally take an equal amount of time for the agent to process (because $H=G$), but only the former has a utility benefit for the agent. A principal who was sending signals about $x_{2}$ could switch to sending signals about $x_{1}$, relax the agent’s participation constraint, and then profit by adjusting the signals so that they induced even more extreme beliefs about $x_{1}$ for the agent.
Proposition 4.
Let $x=(x_{1},x_{2})$, where $x_{1}\in X_{1}$ and $x_{2}\in X_{2}$ and $x_{1},x_{2}$ are independently distributed according to $\bar{q}_{0}$. Suppose (i) $\hat{u}(x)$ depends on $x_{1}$ only and (ii) $G=H$, then (R) is solved by $\pi$ that reveals $x_{1}$ only, i.e. for all $q\in\text{Supp}(\pi)$ and all functions $f:X_{2}\rightarrow\mathcal{R}$, $\mathbb{E}^{q}[f(x_{2})]=\mathbb{E}^{\bar{q}_{0}}[f(x_{2})]$.
Proof.
See Appendix A.5. ∎
Leading Case: Decision-Irrelevant $G$
Let us now consider a different leading case, in which the information that benefits the principal is irrelevant to the agent’s decision. Maintaining the assumption of a product structure for the state space and an independent prior, we now suppose that $G$ depends only on $x_{2}$ (recall that $\hat{u}$ depends only on $x_{1}$. To derive a sharp result, we will assume that the $H$ function is the negative of Shannon’s entropy. [^15] We conjecture that the property required for the result is the $R$-mononticity property discussed in Hébert and La’O (2023).
Proposition 5.
Let $x=(x_{1},x_{2})$, where $x_{1}\in X_{1}$ and $x_{2}\in X_{2}$ and $x_{1},x_{2}$ are independently distributed according to $\bar{q}_{0}$. Suppose (i) $\hat{u}(x)$ depends on $x_{1}$ only, (ii) $G$ depends on $x_{2}$ only and (iii) $H$ is the negative of Shannon’s entropy, then (R) is solved by $\pi=\pi^{*}_{1}\otimes\pi_{2}$, where $\pi^{*}_{1}\in\mathcal{P}(\mathcal{P}(X_{1}))$ is optimal for the agent and $\pi_{2}\in\mathcal{P}(\mathcal{P}(X_{2}))$.
Proof.
See Appendix A.6. ∎
When the principal cares only about providing decision-irrelevant information and the rate of information acquisition is bounded by Shannon’s entropy, the principal will provide a mix of decision-relevant and decision-irrelevant information. The decision-relevant information will be exactly the information that the agent would have chosen for herself, but the agent will be unable to avoid learning the decision-irrelevant information that is “mixed in” with the decision-relevant information.
3.4 Implications of Engagement Maximization
In this subsection, we consider several implications of our results thus far.